# A Midsummer Meme’s Dream: Investigating Market Manipulations in the Meme Coin Ecosystem

### v1

@@ 9. Conclusion and Future Work
Meme coins have rapidly expanded into the cryptocurrency market. However, their community-driven nature makes them highly susceptible to manipulation. This study presents the first large-scale cross-chain analysis of meme coins, covering 34,988 tokens across four major blockchains. We found that 82.6% of high-performing meme coins exhibited clear signs of artificial growth strategies, including wash trading and LPI schemes, and concentrated token ownership. This result suggests that the dramatic gains often associated with meme coins are more likely the result of coordinated manipulation than organic market dynamics. We also detected pump and dumps and strategic rug pulls, with the majority of the tokens involved having previously experienced artificial growth tactics, indicating that early manipulations often serve as precursors to larger exploitation strategies. Finally, we also documented this manipulation pattern in tokens delisted from established aggregators, discovering that 67.74% showed signs of artificial growth manipulation before delisting, and 24.19% exhibited profit extraction tactics after gaining visibility through listing.
@@
These findings underscore the urgent need for stronger investor protection measures in the meme coin ecosystem. Future work should focus on developing real-time detection systems capable of identifying artificial growth tactics such as wash trading and LPI operations, as well as tracking how these evolve into profit extraction schemes. Finally, to gain a deeper understanding of manipulation strategies, it would be valuable to examine the off-chain promotional activities of projects involved in market manipulation. This includes analyzing their websites and social media to understand their strategies and study how online engagement is coordinated with the timing of manipulative activities.
Appendix A Appendix
@@
Figure 7. TF-IDF scores of keywords associated with meme coin names.
[…]
A plot showing the TF-IDF scores of keywords extracted from meme coins’ names.
Figure 8. CDF of the market capitalization of meme coins collected.
[…]
A plot showing the CDF of the market capitalization of meme coins collected.

### v2

@@ 8 Conclusion and Future Work
Meme coins have rapidly expanded into the cryptocurrency market. However, their community-driven nature makes them highly susceptible to manipulation. This study presents the first cross-chain analysis of meme coins across four major blockchains, examining how different manipulation strategies are combined and executed sequentially rather than studying them in isolation. We found that 82.8% of high-performing meme coins exhibited clear signs of artificial growth, including wash trading and LPI, and concentrated token ownership. This result suggests that the dramatic gains often associated with meme coins are more likely the result of coordinated manipulation than organic market dynamics. We also detected pump and dumps and rug pulls, with the majority of the tokens involved having previously experienced artificial growth, indicating that early manipulations often serve as precursors to larger exploitation strategies. Finally, we documented this pattern also in tokens delisted from established aggregators, discovering that 67.74% showed signs of artificial growth before delisting, and 24.19% exhibited profit extraction tactics after gaining visibility through listing. Crucially, we quantified the impact of these exit schemes, linking them to over 17,000 victim addresses and cumulative realized losses exceeding \$9.3 million. These results challenge the perception of meme coins as harmless speculation, revealing instead a structured environment enabling systematic wealth transfer from uninformed to manipulative actors.
@@
This work underscores the urgent need for stronger investor protection measures in the meme coin ecosystem. Future work should focus on developing real-time detection systems capable of identifying wash trading and LPI operations, as well as tracking how these evolve into profit extraction schemes. Finally, to gain a deeper understanding of these strategies, it would be valuable to examine the off-chain promotional activities of projects involved in market manipulation. This includes analyzing their websites and social media to study how online engagement is coordinated with the timing of manipulative activities.
Acknowledgments
This work has been partially funded by MUR National Recovery and Resilience Plan, SERICS (PE00000014).
Ethical Considerations
Data Collection and Privacy.
This study relies exclusively on publicly available data collected from open and accessible sources, including blockchain explorers (e.g., Etherscan, BscScan, Solana RPC), cryptocurrency tracking platforms (e.g., CoinGecko, CoinMarketCap), decentralized exchange aggregators (e.g., DexScreener), and blockchain analytics services (e.g., Dune Analytics). All data pertain to smart contracts, token transactions, and aggregated market information; the study did not involve any interaction with human subjects or the collection of personally identifiable information (PII). Furthermore, throughout our investigation, we made no attempts to correlate blockchain addresses with external off-chain information that could compromise user privacy. In particular, data collection was conducted responsibly. All scraping scripts were designed to respect the operational stability of the target platforms, with requests spaced by at least five seconds to avoid overloading any external service. Additionally, no financial transactions were made during this study, and the authors did not engage in the trading, creation, or promotion of any meme coins examined in this work.
Positive Impacts.
Several stakeholders benefited by this research.
Investors & End-Users. The primary beneficiaries are current and potential investors. Our work provides two key benefits: education and direct protection. It serves as a large-scale, empirical warning, moving beyond anecdotal evidence to show that manipulation is widespread and systemic, especially in high-performing tokens. Additionally, our detection methods are designed to be lightweight and suitable for integration by third-party platforms to provide real-time risk scores, protecting users from investing in tokens that are actively being manipulated.
Information Aggregators. Our findings also provide information aggregators with new, data-driven heuristics to improve their systems. For crypto data aggregators like CoinMarketCap and CoinGecko, which require tokens to meet specific criteria, our research reveals how malicious actors exploit this trust using artificial growth strategies like wash trading and Liquidity Pool-Based Price Inflation (LPI). Our work provides these platforms with detection tools to check tokens more rigorously before listing, thereby maintaining the trust upon which their business models rely. Concerning DEX aggregators like DexScreener and CoinSniper, these platforms list new tokens almost instantly, making them essential for "discovering" new assets. Their problem is twofold: First, their security metrics (e.g., token ownership concentration) are often buried at the bottom of the page and are difficult for non-expert users to understand; second, their warnings are often generic and insufficient (e.g., "this token could be a scam… DYOR"). By adopting our specific detection methods, these platforms could display prominent, actionable warning banners, such as flagging signs of LPI, at the top of token pages. This would allow them to differentiate themselves as safer platforms for discovering new assets.
The S&P Community. We make three key contributions for the community: (1) We present a novel, large-scale dataset focused on meme coins across multiple blockchains, providing a foundation for future studies in this area; (2) We define and provide a detection method for a new manipulation technique (LPI); (3) We highlight the need for future work to study manipulations in combination rather than as isolated events.
Regulators. As regulators begin to deploy new frameworks like MiCA to combat market manipulation (Art. 91 [MiCA2023_Art91]), our work shows how prevalent and how systemic these manipulations are, identifying also novel manipulation techniques (LPI). Our findings are therefore essential for regulators to understand the specific, on-chain mechanisms they will need to detect in order to successfully enforce anti-abuse provisions.
Negative Impacts
Malicious Actors (Manipulators). We acknowledge the dual-use nature of this research. By detailing the mechanisms of novel or complex manipulations like LPI, there is a risk of educating new or less-sophisticated actors about techniques they may not have been aware of. Furthermore, for already-sophisticated manipulators, our paper could be seen as a guide to refining their attacks to evade the specific detection heuristics we propose.
Mitigation and Decision to Publish
The manipulation techniques we describe, particularly LPI, are already in widespread use, as our study demonstrates: Over 17,000 addresses suffered quantifiable harm, including \$3.27M in realized losses from pump-and-dump schemes and \$6.04M in direct losses from rug pulls. This knowledge asymmetry currently favors malicious actors who operate with impunity, while victims and platforms lack detection tools. Public disclosure shifts this balance. Our paper serves to arm the defenders, users, platforms, regulators, and researchers, with the same knowledge, which is the first and most critical step toward building a defense.
By publishing this work, we enable immediate platform protection, evidence-based regulatory action, advanced detection research, and investor education. Thus, we conclude that the clear harm to investors and the defensive benefits of disclosure substantially outweigh the marginal risk of informing actors who largely already possess this knowledge.
Open Science
In adherence to open science principles and to support transparency and reproducibility of our findings, we publicly release on Zenodo (https://doi.org/10.5281/zenodo.17830943) all datasets and results of this study, as well as the methodological implementations employed to detect anomalies, artificial growth operations (wash trading and LPI activities), and profit extraction schemes (rug pulls and pump and dumps).
Specifically, the repository includes the complete dataset of 34,988 confirmed meme coins (Sec.3.1.3), the set of tokens validated across Ethereum, BNB Smart Chain, Solana, and Base (Sec. 3.2.2), and data about creation date of the smart contract (Sec. 3.2.3). It also contains token price data spanning the three-month observation window, which support the price performance analysis presented in Sec. 4.2.
In addition, we provide labeled subsets of tokens used throughout the analyses: high-performing meme coins, delisted tokens, and those exhibiting signs of suspicious behavior, including ownership concentration anomalies, artificial growth operations (e.g., wash trading and Liquidity Pool-Based Price Inflation, or LPI), and profit extraction schemes such as rug pulls and pump-and-dump events (Sec. 5 and Sec. 6).
The dataset is complemented by OHLCV (Open, High, Low, Close, Volume) data at both daily and hourly granularity, which were used to detect manipulations. We also release the transaction-level data and SQL queries used to extract it from Dune Analytics, which we leveraged to validate suspicious activities related to wash trading and LPI (Sec. 5.2.2 and Sec. 5.3.1).
Finally, we release a complete set of Python notebooks implementing the core manipulation detection methods described in the paper, from anomaly identification to the sequential analysis of artificial growth manipulations and profit extraction strategies.
Appendix
@@ Acknowledgments
Figure 7: TF-IDF scores of keywords associated with meme coin names.
Figure 8: CDF of the market capitalization of meme coins collected.